This is from a worksheet from K12-OHVA- My internet school.
How do I find the sum of this finite geometric series? I will use an E for the Sigma notation.
8
E <-1/2)^i-1
i=1

Thank you

Answer by stanbon(75887)

find the sum of this finite geometric series? I will use an E for the Sigma notation.
8
E <-1/2)^i-1
i=1
-------------------
Write out the 8 terms and add them.
------
i = 1 gives you (-1/2)^(1-1) = (-1/2)^0 = 1
i = 2 gives you (-1/2)^(2-1) = (-1/2)^1 = -1/2
i = 3 gives you (-1/2)^(3-1) = (-1/2)^2 = 1/4
i = 4 gives you................(-1/2)^3 = -1/8
i = 5 gives you................(-1/2)^4 = 1/16
i = 6 gives you................(-1/2)^5 = -1/32
i = 7 gives you................(-1/2)^6 = 1/64
i = 8 gives you................(-1/2)^7 = -1/128
-------------------------------------------------------
Adding in pairs you get (1/2) + (1/8) + (1/32) + (1/128) = [64 + 16 + 4 + 1]/128
= 85/128
============================
Cheers,
Stan H.